A rod of length Lo = 1m, moves with speed v equivalent to 77% of the speed of light c along the y-axis direction. At rest, the rod makes an angle of 35 ° with respect to the x- axis.
a) Determine the length of the rod as measured by a stationary observer.
b) Determine the angle θ the rod makes with the x axis when it reaches its speed.
1
Expert's answer
2020-07-06T15:12:54-0400

a). Since, initially angle made with x axis is "35^\circ" thus

"L_{ox}=1\cdot\cos(35^\circ)=0.82m\\
L_{oy}=1\cdot\sin(35^\circ)=0.57m\\"

As there is only motion along y axis, and "v=\frac{77}{100}c"


"L_y=\frac{L_{oy}}{\sqrt{1-v^2/c^2}}\implies L_y=0.89m"

Thus,

"L=\sqrt{L_y^2+L_{ox}^2}=1.21m"

b)Now,

"\theta=\tan^{-1}(L_y/L_{0x})=47.34^\circ"